2. Express in closed form. sum =-infinity to infinity{1/ {z^ 3 -n^ 3}}


1
Expert's answer
2021-10-29T02:53:33-0400

"\displaystyle{\sum_{-\infin}^{\infin}}\frac{1}{z^3-n^3}=2\pi i\cdot \sum res\ f(z_0)" for "z_0" in the upper half-plane


1) "z_0=n"

"res_1\ f(z_0)=\displaystyle{\lim_{z\to n}}[f(z)(z-n)]=\displaystyle{\lim_{z\to n}}\frac{1}{z^2+zn+n^2}=\frac{1}{3n^2}"


2) "z_0=n(-0.5+0.5i\sqrt 3)"

"res_2\ f(z_0)=\displaystyle{\lim_{z\to z_0}}[f(z)(z-z_0)]=\displaystyle{\lim_{z\to z_0}}\frac{1}{z^2+zz_0+z_0^2}=\frac{1}{3z_0^2}=\frac{1}{3n^2(-0.5+0.5i\sqrt 3)^2}"


3) "z_0=n(-0.5-0.5i\sqrt 3)" is not in the upper half-plane

"\displaystyle{\sum_{-\infin}^{\infin}}\frac{1}{z^3-n^3}=\frac{2\pi i}{3n^2}(1+\frac{1}{(-0.5+0.5i\sqrt 3)^2})"